OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through the point (-4, 1) and has a slope of -3/2.

OpenStudy (anonymous):

y-1=-3/2 (x-(-4)) can you finish it?

OpenStudy (anonymous):

No =/

OpenStudy (anonymous):

y-1=-3/2 (x+4) y=-3/2 x -3/2 *4 +1 y=-3/2 x -5

OpenStudy (anonymous):

use this slope point formula y-y1=m(x-x1) y-1=-3/2(x+4) 2(y-1)=-3(x+4) 2y-2=-3x-12 2y+3x-2+12=0 3x+2y+10=0

OpenStudy (anonymous):

need equation in slope-intercept form... it would be: y=-3/2 x -5